# AdaSTaR: Adaptive Data Sampling for Training Self-Taught Reasoners

### v1

@@ 5 Discussion and Additional Empirical Takeaways
We empirically find that the viability of self-improvement via STaR depends on the difficulty of the task relative to the strength of the base model. Therefore, as discussed in § 4.1 and Appendix F, while STaR-based approaches fail to self-improve on Llama 3.2 3B, self-improvement can be realized on the better pre-trained Qwen 2.5 3B. This potentially explains why Peng et al. (2025) uses an instruction-tuned base model instead of cold starting from a pre-trained-only model. Similarly, Hosseini et al. (2024) and Zeng et al. (2025) includes an SFT stage prior to the self-improvement stage. Aligned with recent large reasoning model training (DeepSeek-AI, 2025; Team, 2025; Liu et al., 2025b), the key takeaway is that a STaR-based algorithm is part of a larger training pipeline. Precisely which stage within the training pipeline it should be integrated into is an open problem.
Lastly, we also discuss the limitations and future avenues in Appendix § J and broader impact in Appendix § K.
@@ Appendix A Total Training Time Comparison
Fig. 5 is conducted using full fine-tuning of Llama 3.2 3B (Team, 2024b). The training run-time is set to the early-stopped epoch (iteration) (Caruana et al., 2000).
Figure 5: Total training run-time on 4$\times$×RTX 3090 24G, across three common reasoning datasets CLadder 1.5, ANLI, CQA, SVAMP, GSM8K, and ARC-C. STaR, SFT, and SFT 8-shot Chain-of-Thought is evaluated under zero-shot greedy decoding. Training times across SFT 5-sample Self-Consistency, SFT 8-shot Chain-of-Thought, and SFT differ as their best early-stop epoch differs.
@@ Appendix C AdaD Induces Poor Quality CoT
[…]
@@ D.2 Existing Solution
However, as mentioned in § 2.1, all methods that resolve this excessive sampling ( $M^{t}$-- $\beta^{t}$) problem of STaR (Hosseini et al., 2024; Pang et al., 2024; Zeng et al., 2025; Lin et al., 2025; Peng et al., 2025) simply removes this pre-determined $\beta^{t}$, and instead set $\beta^{t}$ to $|\mathcal{D}^{t}_{+}|$. This approach can be viewed as bringing the blue curve up to the red curve; i.e., $\beta^{t}\leftarrow|\mathcal{D}^{t}_{+}|$. We name this approach as -Full in our experiments (§ 4). B-STaR (Zeng et al., 2025) also embodies this approach. ReSTEM (Singh et al., 2024) does not resolve this problem as they keep the filtering mechanism, as it is useful in their algorithm.
Figure 8: Visualizing the CoT sampling inefficiencies in STaR across numerous datasets. Initial $\beta^{t=1}=40$ as presented in the original implementation. $\beta^{t}$ rises over time as we follow the $+20\%$ of gradient update steps $\sigma$ every iteration in the original implementation. That is, $\beta^{t+1}:=1.2(\beta^{t})$. If $M^{t}$ $>$ $\beta^{t}$, there is an inference sampling inefficiency as $M^{t}$-- $\beta^{t}$ data samples are not used.
@@ Appendix E Further Details on Experimental Configuration and Setting
We primarily conduct our experiments on numerous nodes with 8×8$\times$8 ×RTX 3090 24G, with equivalent hardware specifications across nodes. For a few compute heavy experiments we use nodes with 8×8$\times$8 ×A100 40G. All training is done on the same arbitrary seed value of 10. This value has never been changed. Hyperparameters are organized in Tab. 3.
@@ Appendix G Visualizing Empirical Heaps
[…]
STaR-Full $83.2$ $22$ it $606.2$ $72.6$ $4$ it $684.8$ $91.5$ $16$ it $196.2$
STaR-Acc $84.4$ $11$ it $264.1$ 77.0 $3$ it 305.2 95.0 $10$ it $129.6$
STaR-Acc-Full 84.6 $4$ it 110.8 $74.6$ $2$ it $333.0$ $93.5$ $6$ it 73.0
STaR-Acc-Full-K $82.2$ $2$ it 225.1 77.0 $2$ it $1456.5$ $92.0$ $2$ it $105.3$
ReSTEM $81.0$ $8$ it $874.6$ 77.0 $2$ it $2229.1$ $92.0$ $10$ it $677.2$
B-STaR $83.2$ $10$ it $583.3$ $72.6$ $2$ it $1185.7$ $91.0$ $3$ it $150.4$
AdaSTaR (ours) 85.0 $12$ it $239.9$ ( $\downarrow$ 0%) 77.0 2 it 19.3 ( $\downarrow$ 93.7%) 94.5 $8$ it 83.9 ( $\downarrow$ 35.3%)
@@ Appendix I Gemma 7B Base Model Empirical Results
SFT $49.2$ $0.5$ e $0.01$ $66.0$ $5$ e $1.1$ $61.0$ $3$ e $0.04$
SFT + 8-CoT $76.6$ $5.5$ e $0.13$ 53.0 $5.5$ e $1.2$ $82.0$ $4$ e $0.05$
SFT + 5-SC $66.0$ $0.5$ e $0.01$ $67.6$ $7$ e $1.6$ $61.0$ $3$ e $0.04$
[…]
STaR-Full $76.2$ $3$ it 93.5 $43.8$ $12$ it $5536.6$ $67.5$ $36$ it $832.9$
STaR-Acc 85.4 $13$ it $383.0$ $62.0$ $20$ it $9298.8$ 87.0 $13$ it 281.3
STaR-Acc-Full $84.6$ $20$ it 533.0 $61.8$ $8$ it 3751.0 87.5 $24$ it $510.6$
STaR-Acc-Full-K 85.0 $12$ it $1334.4$ 65.0 $12$ it $20743.3$ 87.5 $6$ it $485.7$
ReSTEM $81.6$ $5$ it $1221.0$ 62.2 17 it 43060.8 $82.5$ $7$ it $1245.6$
@@
AdaSTaR (ours) 85.4 $14$ it 321.0 ( $\downarrow$ 16.2%) 65.2 $20$ it 3055.4 ( $\downarrow$ 85.3%) 89.5 $14$ it 207.0 ( $\downarrow$ 57.4%)
Appendix J Limitation and Future Work
We discuss relevant limitations, to the best of our knowledge, and avenues for future research. First, a natural direction for future work is to explore combinations of AdaSTaR with other advanced STaR-based methods. For instance, investigating the performance of a combined AdaSTaR and an inference-time verifier, such as that in V-STaR (Hosseini et al., 2024), presents a promising research avenue. Such explorations are beyond the scope of the current study. Second, while our experiments demonstrate AdaSTaR’s efficacy, a larger computational budget would have permitted evaluation on even larger-scale models. Nevertheless, our empirical study provides robust evidence of AdaSTaR’s effectiveness across three distinct models: Llama 3.2 3B, Qwen 2.5 3B, and Gemma 7B. Moreover, existing work (Singh et al., 2024) suggests that gains from STaR-based training on smaller models often amplify on larger scales, implying our findings may well extend or even strengthen with increased model size. Furthermore, the model sizes used in our study (up to 7B parameters) are comparable to those in related STaR literature (Zelikman et al., 2022, 2024; Zeng et al., 2025) that uses 6 to 7B base models. Third, similar to other adaptive methods such as Adam (Kingma and Ba, 2015) and AdaGrad (Duchi et al., 2011), AdaSTaR introduces a new hyperparameter $f(\alpha):=\alpha^{2}$. A more granular tuning is deferred to future work. It is anticipated that such tuning could lead to further enhancements in AdaSTaR’s performance and efficiency. Finally, building upon our discussion (§ 5), a salient direction for future work is to investigate the optimal integration of various STaR-based methods within the end-to-end training pipeline incorporating RL-style long CoT generation. This investigation is particularly pertinent given the current divergence in methodologies: the STaR stage is implemented either prior to RL (Team, 2025; Liu et al., 2025b) or subsequent to it (DeepSeek-AI, 2025). Furthermore, a critical open question is whether, and to what extent, enhancements achieved during the STaR phase directly propagate to performance gains in the subsequent RL stage.
Appendix K Broader Impact

### v3

@@ 5 Discussion and Additional Empirical Takeaways
We empirically find that the viability of self-improvement via STaR depends on the difficulty of the task relative to the strength of the base model. Therefore, as discussed in § 4.1 and Appendix F, while STaR-based approaches fail to self-improve on Llama 3.2 3B, self-improvement can be realized on the better pre-trained Qwen 2.5 3B. This potentially explains why Peng et al. (2025) uses an instruction-tuned base model instead of cold starting from a pre-trained-only model. Similarly, Hosseini et al. (2024) and Zeng et al. (2025) includes an SFT stage prior to the self-improvement stage. Aligned with recent large reasoning model training (DeepSeek-AI, 2025; Kimi Team, 2025; Liu et al., 2025b), the key takeaway is that a STaR-based algorithm is part of a larger training pipeline. Precisely which stage within the training pipeline it should be integrated into is an open problem.
6 Limitation and Future Work
We discuss relevant limitations, to the best of our knowledge, and avenues for future research. First, a natural direction for future work is to explore combinations of AdaSTaR with other advanced STaR-based methods. For instance, investigating the performance of a combined AdaSTaR and an inference-time verifier, such as that in V-STaR (Hosseini et al., 2024), presents a promising research avenue. Such explorations are beyond the scope of the current study. Second, while our experiments demonstrate AdaSTaR’s efficacy, a larger computational budget would have permitted evaluation on even larger-scale models. Nevertheless, our empirical study provides robust evidence of AdaSTaR’s effectiveness across three distinct models: Llama 3.2 3B, Qwen 2.5 3B, and Gemma 7B. Moreover, existing work (Singh et al., 2024) suggests that gains from STaR-based training on smaller models often amplify on larger scales, implying our findings may well extend or even strengthen with increased model size. Furthermore, the model sizes used in our study (up to 7B parameters) are comparable to those in related STaR literature (Zelikman et al., 2022, 2024; Zeng et al., 2025) that uses 6 to 7B base models. Third, similar to other adaptive methods such as Adam (Kingma and Ba, 2015) and AdaGrad (Duchi et al., 2011), AdaSTaR introduces a new hyperparameter $f(\alpha):=\alpha^{2}$. A more granular tuning is deferred to future work. It is anticipated that such tuning could lead to further enhancements in AdaSTaR’s performance and efficiency. Finally, building upon our discussion (§ 5), a salient direction for future work is to investigate the optimal integration of various STaR-based methods within the end-to-end training pipeline incorporating RL-style long CoT generation. This investigation is particularly pertinent given the current divergence in methodologies: the STaR stage is implemented either prior to RL (Kimi Team, 2025; Liu et al., 2025b) or subsequent to it (DeepSeek-AI, 2025). Furthermore, a critical open question is whether, and to what extent, enhancements achieved during the STaR phase directly propagate to performance gains in the subsequent RL stage.
Lastly, we discuss broader impact in Appendix § J.
@@ Appendix A Total Training Time Comparison
Fig. 5 is conducted using full fine-tuning of Llama 3.2 3B (Llama Team, 2024). The training run-time is set to the early-stopped epoch (iteration) (Caruana et al., 2000).
[…]
@@ Appendix C AdaD Induces Poor Quality CoT
Correct CoT (STaR-Acc) Given CoT: 6 houses per block, total of 24 pieces per block. $\text{Thus, each house gets: }\frac{24}{6}=4\text{ pieces.}$ Correct Steps: Correctly identified number of houses per block. Correctly distributed mail equally to each house.
@@ D.2 Existing Solution
However, as mentioned in § 2.1, all methods that resolve this excessive sampling ( $M^{t}$- $\beta^{t}$) problem of STaR (Hosseini et al., 2024; Pang et al., 2024; Zeng et al., 2025; Lin et al., 2025; Peng et al., 2025) simply removes this pre-determined $\beta^{t}$, and instead set $\beta^{t}$ to $|\mathcal{D}^{t}_{+}|$. This approach can be viewed as bringing the blue curve up to the red curve; i.e., $\beta^{t}\leftarrow|\mathcal{D}^{t}_{+}|$. We name this approach as -Full in our experiments (§ 4). B-STaR (Zeng et al., 2025) also embodies this approach. ReSTEM (Singh et al., 2024) does not resolve this problem as they keep the filtering mechanism, as it is useful in their algorithm.
Figure 8: Visualizing the CoT sampling inefficiencies in STaR across numerous datasets. Initial $\beta^{t=1}=40$ as presented in the original implementation. $\beta^{t}$ rises over time as we follow the $20\%$ of gradient update steps $\sigma$ every iteration in the original implementation. That is, $\beta^{t+1}:=1.2(\beta^{t})$. If $M^{t}$ >> $\beta^{t}$, there is an inference sampling inefficiency as $M^{t}$- $\beta^{t}$ data samples are not used.
@@ Appendix E Further Details on Experimental Configuration and Setting
We primarily conduct our experiments on numerous nodes with $8\timesRTX$ 3090 24G, with equivalent hardware specifications across nodes. For a few compute heavy experiments we use nodes with $8\timesA$100 40G. All training is done on the same arbitrary seed value of 10. This value has never been changed. Hyperparameters are organized in Tab. 3.
@@ Appendix G Visualizing Empirical Heaps
STaR $80.4$ 2020 it $825.9$ 76.0 44 it $409.2$ $92.5$ 88 it $96.2$
STaR-Full $83.2$ $22$ it $606.2$ $72.6$ 44 it $684.8$ $91.5$ $16$ it $196.2$
STaR-Acc $84.4$ $11$ it $264.1$ 77.0 33 it 305.2 95.0 $10$ it $129.6$
STaR-Acc-Full 84.6 44 it 110.8 $74.6$ 22 it $333.0$ $93.5$ 66 it 73.0
STaR-Acc-Full-K $82.2$ 22 it 225.1 77.0 22 it $1456.5$ $92.0$ 22 it $105.3$
ReSTEM $81.0$ 88 it $874.6$ 77.0 22 it $2229.1$ $92.0$ $10$ it $677.2$
B-STaR $83.2$ $10$ it $583.3$ $72.6$ 22 it $1185.7$ $91.0$ 33 it $150.4$
AdaSTaR (ours) 85.0 $12$ it $239.9$ ( $\downarrow$ 0%) 77.0 2 it 19.3 ( $\downarrow$ 93.7%) 94.5 88 it 83.9 ( $\downarrow$ 35.3%)
@@ Appendix I Gemma 7B Base Model Empirical Results
SFT $49.2$ $0.5$ e $0.01$ $66.0$ 55 e $1.1$ $61.0$ 33 e $0.04$
SFT + 8-CoT $76.6$ $5.5$ e $0.13$ 53.0 $5.5$ e $1.2$ $82.0$ 44 e $0.05$
SFT + 5-SC $66.0$ $0.5$ e $0.01$ $67.6$ 77 e $1.6$ $61.0$ 33 e $0.04$
STaR $82.0$ $17$ it $530.6$ $62.4$ $28$ it $13105.8$ $87.0$ $15$ it $332.8$
STaR-Full $76.2$ 33 it 93.5 $43.8$ $12$ it $5536.6$ $67.5$ $36$ it $832.9$
STaR-Acc 85.4 $13$ it $383.0$ $62.0$ 2020 it $9298.8$ 87.0 $13$ it 281.3
STaR-Acc-Full $84.6$ 2020 it 533.0 $61.8$ 88 it 3751.0 87.5 $24$ it $510.6$
STaR-Acc-Full-K 85.0 $12$ it $1334.4$ 65.0 $12$ it $20743.3$ 87.5 66 it $485.7$
ReSTEM $81.6$ 55 it $1221.0$ 62.2 17 it 43060.8 $82.5$ 77 it $1245.6$
@@
AdaSTaR (ours) 85.4 $14$ it 321.0 ( $\downarrow$ 16.2%) 65.2 2020 it 3055.4 ( $\downarrow$ 85.3%) 89.5 $14$ it 207.0 ( $\downarrow$ 57.4%)
Appendix J Broader Impact